An inductance L and a resistance R are first connected to a battery. After somtime the battery disconnected but L and R remain connected in a closed circuit. Then the current reduces to 37% of its initial value in:


A: RL sec                       B: R/L sec                               C: L/R sec                                 D: 1/LRsec


Grade:10

1 Answers

ROSHAN MUJEEB
3 years ago
When battery disconnected current through the circuit start decreasing exponentially according toi=i0e−Rt/Li=i0e-Rt/L.
⇒0.37i0=i0e−Rt/L⇒0.37=1e=e−Rt/L⇒t=LR